A Stone is thrown horizontally at a speed of 5.0 m/s from the top of a cliff that is 78.4m high. how far from the base of the cliff does the stone hit the ground?
1
Expert's answer
2020-05-03T17:31:18-0400
"x=vt=v\sqrt{\frac{2h}{g}}"

"x=5\sqrt{\frac{2(78.4)}{9.8}}=20\ m"